# Locally Convex Global Loss Network for Decision-Focused Learning

### v2

@@ 5.1 Experimental Settings
Here we explain the experimental details and the baselines used to evaluate our method. Each problem is explained in two stages: the parameter prediction stage and the optimization stage.
Figure 3: A line plot showing the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ for each methodology and problem setting. The metric is lower the better, with 0 representing the optimal value. We tested sample size if $\{2,4,8,16,32\}$ for each problem. The standard error mean for each experiment is detailed in Appendix C.2. Our global surrogate loss LCGLN represented by the red straight line outperforms when 32 samples are used.
@@
We conducted experiments on three different stochastic optimization problems, each presenting unique challenges as explored in previous researches (Donti, Amos, and Kolter 2017; Wilder, Dilkina, and Tambe 2019; Shah et al. 2022). Our experiments were built on top of the public codes from previous researches (Donti, Amos, and Kolter 2017; Shah et al. 2022). Further details on problem formulation can be found in Appendix B.
@@
  • Inventory Stock: The objective of inventory stock problem (Donti, Amos, and Kolter 2017) is to decide the order quantity $a$ to minimize the cost over the stochastic demand $y$. We generate problem instances from randomly sampled $x\in\mathbb{R}^{20}$ with the non-linear mapping $p(y|x;\theta)\propto\text{exp}(\Theta^{\top}x)$ for the fixed true parameter $\Theta\in\mathbb{R}^{20\times 5}$. We assume the demands are discrete with values $d_{1},d_{2},...,d_{5}$ as $1,2,5,10,20$ with probabilities $(p_{\theta})_{i}\equiv p(y=d_{i}|x;\theta)$.
    Note that the realization $y$ of $p_{\theta}$ is known while the true distribution $p_{\theta}$ is unknown; similar to the real-world settings. Since, the true distribution $p_{\theta}$ is unknown, we transform $y$ to a $p_{\theta}$ form when training the surrogate loss functions.
    Prediction: We predict $p_{\theta}$, the distribution of the demand $y$.
    Optimization: Given the demand distribution $p_{\theta}$, we minimize the cost:
    $\min_{a\in\mathbb{R},\quad a_{b},a_{h}\in\mathbb{R}^{k}}$ $c_{0}a+\frac{1}{2}q_{0}a^{2}+\sum_{i=1}^{k}(p_{\theta})_{i}\biggl{(}c_{b}(a_{b})_{i}$
    $+\frac{1}{2}q_{b}(a_{b})_{i}^{2}+c_{h}(a_{h})_{i}+\frac{1}{2}q_{h}(a_{h})_{i}^{2}\biggr{)}$
    s.t. $d-a\boldsymbol{1}\leq a_{b},$
    $a\boldsymbol{1}-d\leq a_{h},$
    $a,a_{h},a_{b}\geq 0.$
@@
  • Budget Allocation: The objective of the submodular optimization problem (Wilder, Dilkina, and Tambe 2019) is to choose $B=2$ websites to advertise based on click-through rates (CTRs) of $U=10$ users among $W=5$ websites. We generate the user features $x=Ay$ multiplying a random matrix $A$ to the CTRs $y$. We conduct experiments with $\{0,5,50,500\}$ fake targets to add difficulty by concatenating fake target size of random CTRs to the original CTRs.
    Prediction: Predict the CTRs $\hat{y}_{w}$ for $W$ websites and $U$ users where user features $x_{u}$ for website $u$ are given.
@@
    Optimization: Choose $B$ websites to advertise that maximizes the expected number of users that click on the advertisement at least once:
    $\max_{a\in\mathbb{R}^{W}}$ $\sum_{u=1}^{U}(1-\prod_{w=1}^{W}a_{w}\cdot\hat{y}_{wu})$
    s.t. $a\in\{0,1\}^{W}$
  • Portfolio Optimization: We choose the portfolio optimization problem to test when good prediction leads to good decisions. We use historical daily stock return data from 2004 to 2017 for 50 stocks (Shah et al. 2022).
@@
    Prediction: Given historical daily stock return data, predict the future stock price $y$.
    Optimization: Given correlation matrix $\Sigma$, decide a portfolio weight $a$ where we want to maximize the expected return for a given level of risk:
    $\max_{a\in\mathbb{R}^{50}}$ $y^{\top}a-\gamma\cdot a^{\top}\Sigma a$
    s.t. $\sum_{i=1}^{50}a_{i}=1$
    where the risk aversion constant $\gamma=0.1$.
@@
Figure 4: A histogram presenting normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) for global surrogate loss models in budget allocation with varying number of fake targets. We use 16 samples for learning loss. The metric is lower the better and 0 when optimal. We test with $\{0,5,50,500\}$ fake targets, noting that the problem becomes more challenging as the number of fake targets increases. Our LCGLN shown in red bars outperforms for most settings.
@@
We compare our global surrogate loss network LCGLN with the following baselines from previous researches. When available, we utilized the hyperparameters specified in previous research for the baselines. In cases where these were not provided, we endeavored to fine-tune to perform their best.
  • PFL: A standard approach that trains predictive model on the input features to produce predictions that closely match the observed parameters. We use the negative log-likelihood (NLL) loss for the inventory stock problem and the mean squared error (MSE) loss for the other problems.
  • DFL: A direct approach that differentiate through the solver to derive gradients for training the predictive model. For inventory stock and portfolio optimization problem, we use differentiable QP solver (Donti, Amos, and Kolter 2017). We use multilinear relaxation (Wilder, Dilkina, and Tambe 2019) for the budget allocation problem.
  • LODL: The local surrogate model (Shah et al. 2022) that learns a surrogate loss locally for each instance, creating total $N$ surrogate loss models. To learn a loss model, LODL generates $K$ number of samples for each $N$ instances and solves for the input of the model. We use directed-quadratic surrogate loss (LODL-DQ) from their paper, where it shows the most promising result among the proposed 4 different loss models.
  • LANCER: The global surrogate model (Zharmagambetov et al. 2024) that uses alternating optimization to learn the predictive model and the global surrogate loss model. The prediction samples are sampled from the predictive model and is used to train the surrogate loss. The predictive model is then trained using the gradients from the updated surrogate loss.
  • EGL: The global surrogate model (Shah et al. 2024) that learns a global surrogate loss with feature-based parameterization and model-based sampling. We use directed-quadratic (EGL-DQ) and weighted-MSE (EGL-WMSE) for the baseline, which showed the best results on budget allocation and portfolio optimization problem respectively.
@@
We use the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ as the evaluation metric, where $\mathcal{R}_{test}$ is a test regret with regret defined in Equation 1. For the budget allocation and portfolio optimization problem which are the maximization task, the regrets are calculated by multiplying the negative sign on the corresponding objective. We calculate the worst case regret $\mathcal{R}_{worst}$ for each problem and derive the normalized test regret. For the inventory stock problem, we assumed the worst-case scenario when the company ordered no stock. In the budget allocation problem, the worst-case scenario was assumed when the advertisements were allocated to the lowest predicted CTRs. For the portfolio optimization problem, we considered the worst-case scenario to be when the entire investment was allocated to the stock with the lowest predicted return. Our metric is 0 when optimal and 1 when worst.
We use a predictive model $\mathcal{M}_{\theta}$ as one hidden layer MLP and a learning rate of 0.001. 500 hidden nodes were used for the portfolio optimization and 10 for the other problems. We mirrored the predictive model exactly for the sampling model $\mathcal{M}_{\xi}$. For LCGLN, we employ a single hidden layer ICNN with two nodes per layer, a learning rate of 0.001 and a softplus activation function. A figure illustrating the LCGLN can be found in Appendix A.1. For each global surrogate loss model employing model-based sampling, we selected the learning rate that demonstrated the best performance from $\{0.01,0.05,0.1,0.5,1\}$. We run 10 experiments for each settings to ensure statistical significance. The experiments were performed on a Ryzen 7 5800X CPU and an RTX 3060 GPU with 64GB of RAM.
@@ 5.2 Results
Table 1 shows the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) for the inventory stock, budget allocation with 500 fake targets and portfolio optimization problems. For the surrogate loss models, we used 32 samples to learn the loss. We categorize the methods into four major training paradigms: two-stage, exact differentiation, local surrogate loss and global surrogate loss.
In the inventory stock problem, LODL-DQ, EGL-WMSE and EGL-DQ suffer from high regret, indicating that they do not provide informative gradients. LANCER and LCGLN perform well compared to others due to their expressiveness, as both use neural networks to learn the loss. For the budget allocation problem, we conduct four experiments with $\{0,5,50,500\}$ fake targets. Detailed results with different fake targets are available in Appendix C.1. For the hardest problem with 500 fake targets, the surrogate models tend to show better results than PFL or DFL methods. In the portfolio optimization problem, good prediction in parameter showed better decisions. This is evidenced by PFL and DFL showing better results compared to surrogate models. EGL-WMSE and LCGLN also demonstrated strong performance in our settings.
To reduce computational cost, it is crucial to train the surrogate loss with small sample sizes. We conducted experiments with varying sample sizes, as shown in the line plot in Figure 3. Note that the normalized test regret is better when lower. Since PFL, DFL and LANCER do not vary with changes in number of sample predictions, they show consistent results across all sample sizes. The surrogate models showed decreased regret as the sample size increased. At a sample size of 32, LCGLN outperformed the provided baselines. For detailed experimental results for each sample size and problem, please refer to Appendix C.2.
We also tested different numbers of fake targets, varying in $\{0,5,50,500\}$, to show how models perform on harder problems. We used 16 samples to train the loss model. Figure 4 presents the normalized test regret with SEM for global models across each experimental setting.
@@ 6 Conclusion
In this paper, we propose Locally Convex Global Loss Network (LCGLN), a global and general surrogate loss for DFL. LCGLN utilizes PCINN as a parametric surrogate to approximate task loss. We use PICNN to guarantee the surrogate to be convex near instances while maintaining a general non-convex structure globally. With LCGLN, users do not need to manually select the appropriate parametric function form for the task loss. Consequently, LCGLN can address general DFL problems with a single surrogate loss, regardless of the amount of observed data. We evaluated our method on three stochastic optimization problems, achieving better decision quality with fewer training samples compared to the state-of-the-art baselines. However, despite the expressive power of LCGLN, a limitation remains: achieving high decision quality requires careful selection of the samples. Thus, our future research will focus on identifying sampling strategies to improve the decision quality for surrogate loss models.
@@ A.1 Architecture Details
Here we show the architecture details of our Locally Convex Global Loss Network (LCGLN). The input is given as the pair of sampled prediction $\hat{y}$ and its corresponding instance $y$.
The bluish boxes represent the vectors that ensure convexity with respect to the input, while the reddish boxes correspond to the vectors that do not guarantee convexity. Yellow boxes denote the hidden layers, where the superscript inside the parentheses indicates the dimension, and the subscript identifies the layer number. For instance, $h_{i}^{(z)}$ is an $i$-th hidden layer with $h_{i}^{(z)}\in\mathbb{R}^{z}$.
We depict linear layers as trapezoids in gray when the dimensions change and as rectangles when they remain constant. The hatched box represents the non-negative linear layer. Details for operators are available in Figure 5.
The left side of the figure shows the architecture from input to the first hidden layer. The right side presents how hidden layers are composed. The upstream of the right side is a flow for reddish vector representations, while the downside is for bluish vectors. For the output layer, one may may simply let $z_{i+1}$ as a output vector and ignore $u_{i+1}$.
Figure 5: Architecture of Locally Convex Global Loss Network. Left figure represents the input to the first hidden layer. The right figure shows the hidden to hidden layers for non-convex vectors (reddish, upstream) and convex vectors (bluish, downstream). Output layer can simply be considered as $z_{i+1}$ without $u_{i+1}$.
Appendix B Experiment Details
@@ B.1 Codes for Paper
Codehttps://anonymous.4open.science/r/lcgln-5E21/README.md
@@ B.2 Details in Inventory Stock Problem
where our objective is to minimize the expectation of cost function:
$\operatorname*{minimize}_{a}\quad\boldsymbol{E}_{y\sim p(y|x;\theta)}\left[f(y,a)\right].$ (6)
By the discrete distribution assumption of demand $y$, we can rewrite the optimization problem as:
$\operatorname*{minimize}_{a\in\mathbb{R},a_{b},a_{h}\in\mathbb{R}^{k}}$ $c_{0}a+\frac{1}{2}q_{0}a^{2}+\sum_{i=1}^{k}(p_{\theta})_{i}\left(c_{b}(a_{b})_{i}+\frac{1}{2}q_{b}(a_{b})_{i}^{2}+c_{h}(a_{h})_{i}+\frac{1}{2}q_{h}(a_{h})_{i}^{2}\right)$ (7)
subject to $d-a\boldsymbol{1}\leq a_{b},\;\,a\boldsymbol{1}-d\leq a_{h},\;\,a,a_{h},a_{b}\geq 0.$
where $(p_{\theta})_{i}\equiv p(y=d_{i}|x;\theta)$ for discrete demands $d_{1},\ldots,d_{k}$.
The parameter $c$ and $q$ each represent the coefficient of linear and quadratic costs, respectively. Specifically, $c_{0},q_{0}$ is related to the costs on the amount of product ordered, $c_{b},q_{b}$ and $c_{h},q_{h}$ is related to the amount of under and over ordered. We use the setting $(c_{0},q_{0},c_{b},q_{b},c_{h},q_{h})$ with $(10,2,30,14,10,2)$.

### v4

@@ 5.1 Experimental Settings
In this section, we explain the experimental details and the baselines used to evaluate our method. Each problem is elaborated in two stages: the parameter prediction stage and the optimization stage.
Figure 3: A line plot showing the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ for each methodology and problem setting. The metric is lower the better, with 0 representing the optimal value. We tested sample size of $\{2,4,8,16,32\}$ for each problem. The standard error mean for each experiment is detailed in Appendix C.2. Our global surrogate loss LCGLN represented by the red straight line outperforms when 32 samples are used.
@@
We conducted experiments on three different stochastic optimization problems, each presenting unique challenges as explored in previous research (Donti, Amos, and Kolter 2017; Wilder, Dilkina, and Tambe 2019; Shah et al. 2022). Our experiments were built on top of the public codes from previous research (Donti, Amos, and Kolter 2017; Shah et al. 2022). Here we briefly describe each problem. For further problem descriptions, please refer to Appendix B.
@@
  • Inventory Stock (Donti, Amos, and Kolter 2017): We decide the order quantity to minimize the cost over the stochastic demand. To simplify the problem, we assume the demands are discrete.
    Prediction: We predict the discrete probability distribution of the demand.
    Optimization: With the predicted demand distribution, we decide the order quantity that minimizes the cost.
@@
  • Budget Allocation (Wilder, Dilkina, and Tambe 2019): We choose websites to advertise based on click-through rates (CTRs) of users among websites. As a variant, we conduct four different experiment settings by concatenating $\{0,5,50,500\}$ size of random CTRs (fake targets) to the original CTRs and adjusting the problem difficulty.
    Prediction: Given the website features, we predict the CTRs.
@@
    Optimization: With the predicted CTRs, we choose the websites to advertise that maximize the expected number of users who click on the advertisement at least once.
  • Portfolio Optimization (Shah et al. 2022): We allocate weights for invested stocks to maximize the expected risk-adjusted return.
@@
    Prediction: Given the historical daily stock return data, we predict the future stock price.
    Optimization: Using the predicted stock price, we assign portfolio weights that maximize the expected risk-adjusted return, given the covariance matrix.
@@
Figure 4: A histogram presenting normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) for global surrogate loss models in budget allocation with varying number of fake targets. We use 16 samples for learning loss. The metric is lower the better and 0 when optimal. We test with $\{0,5,50,500\}$ fake targets, noting that the problem becomes more challenging as the number of fake targets increases. Our LCGLN shown in red bars outperforms most settings.
@@
We compare our global surrogate loss network LCGLN with the following baselines from previous research. When available, we utilized the hyperparameters specified in previous research for the baselines. In cases where these were not provided, we endeavored to fine-tune them to perform their best.
  • PFL: A standard approach that trains a predictive model on the input features to produce predictions that closely match the observed parameters. We use the negative log-likelihood (NLL) loss for the inventory stock problem and the mean squared error (MSE) loss for the other problems.
  • DFL: A direct approach that differentiates through the solver to derive gradients for training the predictive model. For the inventory stock and the portfolio optimization problems, we use the differentiable QP solver (Donti, Amos, and Kolter 2017). We use multilinear relaxation (Wilder, Dilkina, and Tambe 2019) for the budget allocation problem.
  • LODL: The local surrogate model (Shah et al. 2022) that learns a surrogate loss locally for each instance, creating total $N$ surrogate loss models. To learn a loss model, LODL generates $K$ number of samples for each $N$ instance and solves for the input of the model. We use directed-quadratic surrogate loss (LODL-DQ) from their paper, the most promising result among the proposed 4 different loss models.
  • LANCER: The global surrogate model (Zharmagambetov et al. 2024) that uses alternating optimization to learn the predictive model and the global surrogate loss model. The prediction samples are sampled from the predictive model and are used to train the surrogate loss. The predictive model is then trained using the gradients from the updated surrogate loss.
  • EGL: The global surrogate model (Shah et al. 2024) that learns a global surrogate loss with feature-based parameterization and model-based sampling. We use directed-quadratic (EGL-DQ) and weighted-MSE (EGL-WMSE) for the baseline, which showed the best results on the budget allocation and the portfolio optimization problems respectively.
@@
We use the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ as the evaluation metric, where $\mathcal{R}_{test}$ is a test regret with regret defined in Equation 1. For the maximization tasks, such as the budget allocation and portfolio optimization problem, regrets are calculated by multiplying the negative sign on the corresponding objective. We calculate the worst case regret $\mathcal{R}_{worst}$ for each problem and derive the normalized test regret. For the inventory stock problem, we assumed the worst-case scenario when the company ordered no stock. In the budget allocation problem, the worst-case scenario was assumed when the advertisements were allocated to the lowest predicted CTRs. For the portfolio optimization problem, we considered the worst-case scenario to be when the entire investment was allocated to the stock with the lowest predicted return. Our metric is 0 when optimal and 1 when worst.
We used a predictive model $\mathcal{M}_{\theta}$ as one hidden layer MLP and a learning rate of 0.001. 500 hidden nodes were used for the portfolio optimization and 10 for the other problems. We mirrored the predictive model exactly for the sampling model $\mathcal{M}_{\xi}$. For LCGLN, we employed a single hidden layer PICNN with two nodes per layer, a learning rate of 0.001, and a softplus activation function. A figure illustrating the LCGLN can be found in Appendix A.1. For each global surrogate loss model employing model-based sampling, we selected the learning rate that demonstrated the best performance from $\{0.01,0.05,0.1,0.5,1\}$. We run 10 experiments for each setting to ensure statistical significance. The experiments were performed on a Ryzen 7 5800X CPU and an RTX 3060 GPU with 64GB of RAM.
@@ 5.2 Results
Table 1 shows the normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) for the inventory stock, budget allocation with 500 fake targets and portfolio optimization problems. For the surrogate loss models, we used 32 samples to learn the loss. We categorize the methods into four major training paradigms: two-stage, exact differentiation, local surrogate loss, and global surrogate loss.
In the inventory stock problem, LODL-DQ, EGL-WMSE, and EGL-DQ suffer from high regret, indicating that they do not provide informative gradients. LANCER and LCGLN perform well compared to others due to their expressiveness, as both use neural networks to learn the loss. For the budget allocation problem, we conduct four experiments with $\{0,5,50,500\}$ fake targets. Detailed results with different fake targets are available in Appendix C.1. For the hardest problem with 500 fake targets, the surrogate models tend to show better results than PFL or DFL methods. In the portfolio optimization problem, good prediction in parameters showed better decisions. This is evidenced by PFL and DFL showing better results compared to surrogate models. EGL-WMSE and LCGLN also demonstrated strong performance in our settings.
To reduce the computational cost, it is crucial to train the surrogate loss with small sample sizes. We conducted experiments with varying sample sizes, as shown in the line plot in Figure 3. Note that the normalized test regret is better when lower. Since PFL, DFL, and LANCER do not vary with changes in a number of sample predictions, they show consistent results across all sample sizes. The surrogate models showed decreased regret as the sample size increased. At a sample size of 32, LCGLN outperformed the provided baselines. For detailed experimental results for each sample size and problem, please refer to Appendix C.2.
We also tested different numbers of fake targets, varying in $\{0,5,50,500\}$, to show how models perform on harder problems. We used 16 samples to train the loss model. Figure 4 presents the normalized test regret with SEM for global models across each experimental setting. LCGLN consistently outperformed in most settings.
@@ 6 Conclusion
In this paper, we propose Locally Convex Global Loss Network (LCGLN), a global and general surrogate loss for DFL. LCGLN utilizes PICNN as a parametric surrogate to approximate task loss. We use PICNN to guarantee the surrogate to be convex near instances while maintaining a general non-convex structure globally. With LCGLN, users do not need to manually select the appropriate parametric function form for the task loss. Consequently, LCGLN can address general DFL problems with a single surrogate loss, regardless of the amount of observed data. We evaluated our method on three stochastic optimization problems, achieving better decision quality with fewer training samples compared to the state-of-the-art baselines. However, despite the expressive power of LCGLN, a limitation remains: achieving high decision quality requires careful selection of the samples. Thus, our future research will focus on identifying sampling strategies to improve the decision quality for surrogate loss models.
Acknowledgements
This research was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (NRF-2022M3J6A1063021 and NRF-RS-2023-00208980).
@@ A.1 Architecture Details
Here we show the architecture details of our Locally Convex Global Loss Network (LCGLN). The input is given as the pair of the sampled prediction $\hat{y}$ and its corresponding instance $y$.
The bluish boxes represent the vectors that ensure convexity with respect to the input, while the reddish boxes correspond to the vectors that do not guarantee convexity. The yellow boxes denote the hidden layers, where the superscript inside the parentheses indicates the dimension, and the subscript identifies the layer number. For instance, $h_{i}^{(z)}$ is an $i$-th hidden layer with $h_{i}^{(z)}\in\mathbb{R}^{z}$.
We depict linear layers as trapezoids in gray when the dimensions change and as rectangles when they remain constant. The hatched box represents the linear layer with non-negative weights. Details for operators are available in Figure 5.
The left side of the figure shows the architecture from the input to the first hidden layer. The right side presents how hidden layers are composed. The upstream of the right side is a flow for reddish vector representations, while the downside is for bluish vectors. For the output layer, one may simply let $z_{i+1}$ be the output vector and ignore $u_{i+1}$.
Figure 5: Architecture of Locally Convex Global Loss Network. The left figure represents the input to the first hidden layer. The right figure shows the hidden to hidden layers for non-convex vectors (reddish, upstream) and convex vectors (bluish, downstream). The output layer can simply be considered as $z_{i+1}$ without $u_{i+1}$.
Appendix B Experimental Details
@@ B.1 Codes for Paper
Codehttps://github.com/HESEDaniel/lcgln
@@ B.2 Details in Inventory Stock Problem
In the inventory stock problem (Donti, Amos, and Kolter 2017), we consider a situation where one must order $a$ number of products before the demand $y$ is realized to minimize the cost. We generate problem instances from randomly sampled $x\in\mathbb{R}^{20}$ with the non-linear mapping $p(y|x;\theta)\propto\text{exp}(\Theta^{\top}x)$ for the fixed true parameter $\Theta\in\mathbb{R}^{20\times 5}$. To simplify the problem, we assume the demands are discrete with values $d_{1},d_{2},...,d_{k}$ with probabilities $(p_{\theta})_{i}\equiv p(y=d_{i}|x;\theta)$. Note that the realization $y$ of $p_{\theta}$ is known while the true distribution $p_{\theta}$ is unknown. Since, the true distribution $p_{\theta}$ is unknown, we transform $y$ to a $p_{\theta}$ form when training the surrogate loss functions.
@@
where our objective is to minimize the expected cost:
$\min_{a}\quad\boldsymbol{E}_{y\sim p(y|x;\theta)}\left[f(y,a)\right].$ (6)
Here, we let $k=5$ and assign $d_{1},\ldots,d_{5}$ as 1, 2, 5, 10, 20 with probability $(p_{\theta})_{i}\equiv p(y=d_{i}|x;\theta)$ for each discrete demand; identical to the previous research settings (Donti, Amos, and Kolter 2017). Then, we can rewrite the optimization problem as:
$\min_{a\in\mathbb{R},\quad a_{b},a_{h}\in\mathbb{R}^{k}}$ $c_{0}a+\frac{1}{2}q_{0}a^{2}+\sum_{i=1}^{k}(p_{\theta})_{i}\left(c_{b}(a_{b})_{i}+\frac{1}{2}q_{b}(a_{b})_{i}^{2}+c_{h}(a_{h})_{i}+\frac{1}{2}q_{h}(a_{h})_{i}^{2}\right)$ (7)
s.t. $d-a\boldsymbol{1}\leq a_{b},\;\,a\boldsymbol{1}-d\leq a_{h},\;\,a,a_{h},a_{b}\geq 0.$
The parameters $c$ and $q$ each represent the coefficient of linear and quadratic costs, respectively. Specifically, $c_{0}$ and $q_{0}$ are related to the costs of the amount of product ordered. $c_{b},q_{b}$ and $c_{h},q_{h}$ are related to the amount of under and over-ordered. We use the setting $(c_{0},q_{0},c_{b},q_{b},c_{h},q_{h})$ with $(10,2,30,14,10,2)$.
B.3 Details in Budget Allocation Problem
The objective of the submodular optimization problem (Wilder, Dilkina, and Tambe 2019) is to choose $B=2$ websites to advertise based on click-through rates (CTRs) of $U=10$ users among $W=5$ websites. We generate the website features $x_{w}\in\mathbb{R}^{U}$ for each website $w$ by multiplying a random matrix $A\in\mathbb{R}^{U\times U}$ to the CTRs $y_{w}\in\mathbb{R}^{U}$. We conduct experiments with $\{0,5,50,500\}$ fake targets to add difficulty by concatenating the fake target size of random CTRs to the original CTRs. Given the website features $x_{w}$ for each website $w$, we first predict the CTRs $\hat{y}_{w}$. Then, we choose $B$ websites to advertise using a binary vector $a\in\mathbb{R}^{W}$ that maximizes the expected number of users that click on the advertisement at least once:
$\max_{a}$ $\sum_{u=1}^{U}(1-\prod_{w=1}^{W}(1-a_{w}\cdot\hat{y}_{wu}))$
s.t. $a\in\{0,1\}^{W}$
B.4 Details in Portfolio Optimization Problem
For the portfolio optimization problem, we use historical daily stock return data from 2004 to 2017 for $N=50$ stocks (Shah et al. 2022). Using this historical daily stock return data, we predict the future stock price $y\in\mathbb{R}^{N}$. Subsequently, we solve for the portfolio weight $a\in\mathbb{R}^{N}$, aiming to maximize the expected return given the covariance matrix $\Sigma$ and the risk aversion coefficient $\gamma$:
$\max_{a}$ $y^{\top}a-\gamma\cdot a^{\top}\Sigma a$
s.t. $\sum_{i=1}^{N}a_{i}=1$
where we set $\gamma=0.1$.